In Exercises 17-32, a vector u in R and a subspace W of R are given. (a) Find the orthogonal projection matrix Pw. (b) Use19. u = 2 and W is the solution set of the system of . equations X1 + X2 - X3 = 0 x1 - x2 + 3x3 = 0.
Answer #1

19. x1 + x₂ - Xz = 0 () 21-Xy + 3x3 = 0 (2) (1)+ (2) 2x + 2 X3 20 4 = - X3 from (1) - Xz 4 X₂ - X₃ = 0 Xg 2x3 The solutions a